Solve the linear system
2 x + 2y = ,4 3x − y = …………….. (I)
by substitution.
If 3 4x + y = ax,c + by = is a system having the same solution set as (I),
find c,b,a .
1
Expert's answer
2019-03-26T12:32:45-0400

Solve the linear system

2 x + 2y =4

3x − y =4

 

4x + y = ax

c + by =3


In the first system we substitute x from the first equation in the second

x=2-y

3x-y=4


3*(2-y)-y=4


6-4y=4


y=(6-4)/4=0.5


x=2-0.5=1.5



since the solutions of the first system are the same as the solutions of the second, we substitute x and y for the number.


4*1.5+0.5=a1.5

c+0.5b=3


a=4.5/1.5=3

c=3-0.5b

b=6-2c


the values ​​b and c have an infinite number of solutions